OpenStudy (anonymous):

Let g(x)=2x^2+11x. Find a solution that g(a)=40

OpenStudy (campbell_st):

well substitute x with a and f(x) with 40 so you get \[40 = 2a^2 + 11a\] rewrite the equation \[2a^2 + 11a - 40 = 0\] you need to solve the equation, it can be factored. hope this helps.

OpenStudy (nurali):

g(x)=2x^2+11x g(a)=2a^2+11a g(a)=40 40=2a^2+11a 2a^2+11a-40=0 2a^2 + 11a - 40 = 0 2a^2+16a-5a-40=0 2a(a+8)-5(a+8)=0 (2a - 5)(a + 8) = 0 2a-5=0 or a+8=0 2a=5 or a=-8 a=5/2 or a=-8